# FaceCoT: A Benchmark Dataset for Face Anti-Spoofing with Chain-of-Thought Reasoning

### v1

@@ Abstract
Face Anti-Spoofing (FAS) typically depends on a single visual modality when defending against presentation attacks such as print attacks, screen replays, and 3D masks, resulting in limited generalization across devices, environments, and attack types. Meanwhile, Multimodal Large Language Models (MLLMs) have recently achieved breakthroughs in image–text understanding and semantic reasoning, suggesting that integrating visual and linguistic co-inference into FAS can substantially improve both robustness and interpretability. However, the lack of a high-quality vision–language multimodal dataset has been a critical bottleneck. To address this, we introduce FaceCoT (Face Chain-of-Thought), the first large-scale Visual Question Answering (VQA) dataset tailored for FAS. FaceCoT covers 14 spoofing attack types and enriches model learning with high-quality CoT VQA annotations. Meanwhile, we develop a caption model refined via reinforcement learning to expand the dataset and enhance annotation quality. Furthermore, we introduce a CoT-Enhanced Progressive Learning (CEPL) strategy to better leverage the CoT data and boost model performance on FAS tasks. Extensive experiments demonstrate that models trained with FaceCoT and CEPL outperform state-of-the-art methods on multiple benchmark datasets.
@@ 1 Introduction
Face Anti-Spoofing (FAS) plays a vital role in securing face recognition systems, yet it must contend with a wide spectrum of sophisticated presentation attacks such as printed photos, screen-based replay, and 3D masks. The diversity of attack types poses significant challenges to FAS models. However, most existing approaches DA4 ; DA5 ; deep4 ; DG5 ; deep3 ; deep2 rely solely on a single visual modality, which severely limits their generalization across devices, environments, and attack types, and offers no explicit rationale for their decisions, resulting in poor interpretability.
Meanwhile, Multimodal Large Language Models (MLLMs) gpt4o ; internvl-2.5 ; Qwen2.5-VL ; minicpm have achieved remarkable progress in tasks such as image and text understanding, visual question answering and language reasoning. These models can integrate visual and language information to perform causal reasoning and semantic interpretation, making them particularly suited to FAS tasks. Consequently, a multimodal approach of images and text represents a novel solution path for FAS. Through large-scale pre-training, such models have the potential to overcome the limited generalization of existing FAS methods. Moreover, their strong visual-to-text alignment capabilities combined with powerful language reasoning modules can offer clear decision rationale, thereby significantly enhancing the interpretability of model predictions.
However, this paradigm is hindered by the lack of high-quality image and text multimodal datasets for FAS. Available public FAS datasets casia-mfsd ; CelebA-Spoof ; surf-3d-mask ; wild-fas ; replay-attack ; msu-mfsd provide only image or video inputs with binary real or fake labels and omit the structured language information needed for MLLMs training. Using these limited datasets directly for MLLMs training can lead to overfitting and fail to supply an explicit reasoning chain to enhance interpretability. To address this gap, we construct a Chain-of-Thought (CoT) annotated FAS dataset, named FaceCoT.
@@
The construction of FaceCoT proceeds in three stages. First, to guide the model to perform human-like reasoning—inspecting from global context down to local details and deriving a final judgment, we design a dedicated CoT annotation format tailored to FAS. As shown in Figure 1, this format comprises six hierarchical levels: caption, facial description, facial attributes, reasoning, spoofing description, and conclusion. Second, we employ GPT-4o gpt4o to automatically generate initial CoT annotations. To guarantee data quality and accuracy, we adopt a human-in-the-loop workflow in which all automatically generated content is rigorously filtered and refined by expert annotators. The resulting high-quality subset is named FaceCoT-Gold100K. Third, to expand the dataset’s scale, we train a specialized FAS caption model on FaceCoT-Gold100K. During this process, we integrate a rule-verifiable Reinforcement Learning (RL) strategy to improve annotation quality and robustness across domains. The optimized model is capable of producing high-quality CoT annotations even on unseen data, allowing us to generate an additional 982K structured annotations, termed FaceCoT-Silver982K.
@@
To the best of our knowledge, FaceCoT is the first Visual Question Answering (VQA) dataset specifically designed for FAS, aggregating 1.08M training samples from two large-scale public FAS datasets CelebA-Spoof ; wild-fas and covering 14 distinct attack types. The types of attacks included in the dataset are illustrated in Figure 2(a). Its carefully engineered, structured and hierarchical CoT annotation format not only ensures logical consistency and interpretability but also provides a natural learning pathway for downstream reasoning models. Moreover, a hybrid annotation workflow—combining GPT-4o–driven automatic generation with expert manual refinement—further guarantees data quality and reliability. Finally, the integration of an RL–based strategy during the dataset expansion phase enhances cross-domain annotation accuracy and robustness, supplying the multimodal FAS community with unprecedentedly large-scale, high-quality training resources.
@@
The FaceCoT dataset provides a unique resource for training MLLMs. However, an end-to-end scheme forces the model to learn CoT reasoning and classification simultaneously, causing task interference and preventing the reasoning objective from fully converging. As a result, the model cannot fully exploit the fine-grained visual cues embedded in the CoT annotations. To address this, we propose a CoT-Enhanced Progressive Learning (CEPL) strategy, consisting of two stages: (1) Visual Enhancement Pre-training, we perform full-parameter Supervised Fine-Tuning (SFT) of the model using CoT data, thereby focusing the vision encoder on extracting fine-grained, spoof-relevant facial features; (2) Multi-task Joint Training, we inherit the vision encoder from the first stage, reset the connector and language decoder to their original pretrained weights, and insert lightweight adapters into both the connector and decoder. We then jointly train on CoT annotations and binary labels with a multi-task loss, ensuring that the model retains deep reasoning capabilities while rapidly adapting to the classification task.
@@
  • Proposing and releasing the FaceCoT dataset To address the lack of VQA data in multimodal FAS, we introduce FaceCoT—the first VQA dataset designed specifically for FAS. It covers 14 attack types and comprises 1.08M samples. Each entry contains a structured visual CoT question-answering, guiding models from image understanding to logical judgment and greatly enhancing detection accuracy and generalization.
  • Proposing a CoT-Enhanced Progressive Learning method To fully leverage CoT data for FAS detection, we develop a CoT-Enhanced Progressive Learning approach that balances CoT reasoning and binary classification. This strategy significantly improves model performance on both reasoning and classification tasks.
  • Performance Extensive experiments on FAS benchmarks demonstrate the value of our dataset and the effectiveness of our method. As shown in Figure 2(b), our approach outperforms state-of-the-art methods, achieving an average AUC improvement of 5.00% and an HTER reduction of 4.06%.
@@ 2 Related work
FAS is a key technology for protecting face recognition face-recognize systems from presentation attacks. FAS focuses on detecting these fraudulent attacks to ensure that the faces recognized by the system are genuine live faces. Early FAS methods primarily rely on handcrafted low-level feature extraction techniques, such as LBP lbp-top ; lbp , SIFT SIFT , HoG HoG , and others. With advancements in deep learning, many FAS methods deep1 ; deep2 ; deep3 ; deep4 ; deep5 begin using models like CNNs and ViT to train classifiers. FAS methods multi1 ; multi2 ; multi3 gradually introduce multimodal learning, combining features from different types of data (e.g., RGB, IR, depth) to improve model performance. These methods achieve good results in intra-dataset scenarios but struggle with generalization to unseen attack types from out-of-domain data. As a result, Domain Adaptation (DA) DA1 ; DA2 ; DA3 ; DA4 ; DA5 and Domain Generalization (DG) DA1 ; DG2 ; DG3 ; deep4 ; DG5 ; DG6 approaches are developed. DA minimizes the distribution gap between the source and target domains by utilizing unlabeled target data. However, in practice, collecting unlabeled target data is challenging. DG learns broadly applicable features from multiple source domains, enabling good predictions in the target domain. However, due to the diversity of attack types and data collection methods, it is difficult to find a universal feature space for fake faces, leading to insufficient generalization. Additionally, these methods lack interpretability, making it challenging to understand the decision-making process of the model, which is crucial in high-stakes applications like security systems.
@@
In MLLMs, CoT is a method for solving complex problems through a series of intermediate reasoning steps, providing an interpretable reasoning path for the model when answering questions. Recent research cot1 ; cot2 ; cot3 has shown that the CoT prompting method significantly enhances the reasoning abilities of large language models in reasoning tasks. Consequently, researchers have attempted to apply CoT in MLLMs. Some researchers define their own reasoning stages. For example, LLaVA-CoT llavacot proposes a method combining CoT with MLLMs by designing a "summary-caption-reasoning-conclusion" reasoning process to enhance the model’s reasoning ability. Other researchers have explored allowing models to autonomously design the reasoning stages, such as PS-CoT ps-cot , which enables large language models to generate task-solving plans before generating reasoning evidence. CoT-PT cot-pt adopts a hierarchical reasoning approach, moving from abstract concepts to specific ones. BDoG BDoG employs a unique approach using three agents who repeatedly debate to implicitly form a reasoning graph that explores and aggregates various thoughts. However, in the field of FAS, such advanced CoT-based reasoning strategies have not yet been fully explored or applied.
@@ 3 The FaceCoT Dataset
In this section, we introduce the FaceCoT dataset, comprising FaceCoT-Gold100K and FaceCoT-Silver982K, and detail the entire construction process. Figure 3 illustrates the entire data construction pipeline. Subsection 3.1 describes how we collecte and curate our raw data to form FaceCoT-Gold100K. Subsection 3.2 provides a detailed explanation of how we design the CoT structure for FAS and use GPT-4o gpt4o for data annotation. Subsection 3.3 presents the training of our FAS caption model—using FaceCoT-Gold100K under an SFT augmented by RL, and how we then apply it to annotate the complete raw dataset, yielding the larger FaceCoT-Silver982K dataset.
@@ 3.1 Data collection
Our original dataset of FaceCoT-Gold100K is collected from two major FAS datasets: CelebA-Spoof CelebA-Spoof and Wild-FAS (WFAS) wild-fas . CelebA-Spoof is a comprehensive FAS dataset containing 625K images from 10K subjects, including genuine faces and 10 types of spoofing attacks. WFAS is a large, diverse FAS dataset collected in unconstrained environments. It contains 853K images from 321K spoofed subjects and 529K images from 148K genuine subjects, covering 14 types of presentation attacks, including both 2D and 3D forms.
@@ 3.2 Data annotation
Humans typically identify spoofed images by following a hierarchical “global-to-local” reasoning path: they first assess the overall scene, then focus on facial details and attributes, and finally draw a conclusion through logical analysis. To emulate this process, we partition our CoT annotations into six modules, enabling the model to perform human-like spoof detection reasoning:
@@
    Facial Description: This part focuses the model’s attention on the facial region of the image and requires a description of the facial features. The face is the most easily simulated area in spoofing attacks, so this section helps the model concentrate on potential spoofing regions.
@@
    Facial Attributes: Further, the model is asked to describe various facial attributes, including facial features, textures, and expressions. By describing these attributes, the model’s ability to perceive fine-grained details is enhanced.
@@
    Conclusion: The conclusion section is the final summary of all the reasoning conducted previously.
Based on the above design approach, we further standardize the annotation format of FaceCoT data, adopting the following structure: <Caption></Caption>, <Facial Description></Facial Description>, <Facial Attributes></Facial Attributes>, <Reasoning></Reasoning>, <Spoofing Description></Spoofing Description>, <Conclusion></Conclusion>, thus providing the model with clear and structured input, which helps improve the accuracy and stability of multimodal reasoning (as shown in Figure 1).
@@
After designing the CoT format for the answers, we begin the annotation process by calling GPT-4o gpt4o . During annotation, for different spoofing types, we provide corresponding hints (e.g., classification standards such as "photographing a newspaper is considered a spoofing action") to prevent the model from identifying features but failing to properly control the decision boundaries.
@@
After all data annotation, we compare the labels obtained from the answers using regular expressions with the original labels. There are 98,976 correct results, and any incorrect data is re-annotated. Samples that still fail this check are flagged as hard cases, yielding 581 such instances. For these head cases, professional annotators perform cleaning. The annotators review the original labels, locate the corresponding stages, and correct any deficiencies. Finally, they reassemble a logically coherent CoT text. The detailed hard case cleaning process is provided in the Appendix.
@@ 3.3 Data expansion
Although we assembled a diverse FaceCoT-Gold100K, it still falls short of covering every attack type in real-world FAS scenarios. Expanding the dataset with GPT-4o plus human annotation would be costly and hard to scale. Therefore, we train a caption model to address this issue. We first train a caption model on FaceCoT-Gold100K via SFT, but find that its auto-generated CoT annotations on unseen images suffer from two issues: semantic errors and format errors. To fix this, we adopt a verified reinforcement fine-tuning scheme inspired by VRFT VRFT , designing rewards for both accuracy and format compliance. Specifically, an accuracy reward gives 1 if <Conclusion> matches the image’s ground-truth label and 0 otherwise, while a format reward checks whether the output follows the FaceCoT template. This RL step markedly improves annotation correctness and offers a scalable path to enlarging FaceCoT. The details of the RL procedure are provided in the Appendix.
@@
We apply the trained caption model to annotate the training splits of CelebA-Spoof CelebA-Spoof and WFAS wild-fas . Annotation accuracy is measured by (1) regular-expression checks for template compliance and (2) consistency between the generated <Conclusion> tag and the ground-truth label. With standard SFT, we achieve an accuracy of 88%, whereas reinforcement-fine-tuning raises this figure to 99.6%. We ultimately create FaceCoT-Silver982K, an additional 982K high-quality face CoT annotations.

### v2

# Harnessing Chain-of-Thought Reasoning in Multimodal Large Language Models for Face Anti-Spoofing

@@ Abstract
Face Anti-Spoofing (FAS) typically depends on a single visual modality when defending against presentation attacks such as print attacks, screen replays, and 3D masks, resulting in limited generalization across devices, environments, and attack types. Meanwhile, Multimodal Large Language Models (MLLMs) have recently achieved breakthroughs in image–text understanding and semantic reasoning, suggesting that integrating visual and linguistic co-inference into FAS can substantially improve both robustness and interpretability. However, the lack of a high-quality vision–language multimodal dataset has been a critical bottleneck. To address this, we introduce FaceCoT (Face Chain-of-Thought), the first large-scale Visual Question Answering (VQA) dataset tailored for FAS. FaceCoT covers 14 spoofing attack types and enriches model learning with high-quality CoT VQA annotations. Meanwhile, we develop a caption model refined via Reinforcement Learning (RL) to expand the dataset and enhance annotation quality. Furthermore, we introduce a CoT-Enhanced Progressive Learning (CEPL) strategy to better leverage the CoT data and boost model performance on FAS tasks. Extensive experiments demonstrate that models trained with FaceCoT and CEPL outperform state-of-the-art methods on multiple benchmark datasets.
@@ 1 Introduction
Face Anti-Spoofing (FAS) plays a vital role in securing face recognition systems, yet it must contend with a wide spectrum of sophisticated presentation attacks such as printed photos, screen-based replay, and 3D masks. The diversity of attack types poses significant challenges to FAS models. However, most existing approaches [50, 18, 32, 15] rely solely on a single visual modality, which severely limits their generalization across devices, environments, and attack types, and offers no explicit rationale for their decisions, resulting in poor interpretability.
Meanwhile, Multimodal Large Language Models (MLLMs) [5, 1] have achieved remarkable progress in tasks such as image and text understanding, visual question answering and language reasoning. These models can integrate visual and language information to perform causal reasoning and semantic interpretation, making them particularly suited to FAS tasks. Consequently, a multimodal approach of images and text represents a novel solution path for FAS. Through large-scale pre-training, such models have the potential to overcome the limited generalization of existing FAS methods. Moreover, their strong visual-to-text alignment capabilities combined with powerful language reasoning modules can offer clear decision rationale, thereby significantly enhancing the interpretability of model predictions.
However, this paradigm is hindered by the lack of high-quality image and text multimodal datasets for FAS. Available public FAS datasets [12, 3, 42] provide only image or video inputs with binary real or fake labels and omit the structured language information needed for MLLMs training. Using these limited datasets directly for MLLMs training can lead to overfitting and fail to supply an explicit reasoning chain to enhance interpretability. To address this gap, we construct a Chain-of-Thought (CoT) annotated FAS dataset, named FaceCoT.
@@
The construction of FaceCoT proceeds in three stages. First, to guide the model to perform human-like reasoning, which involves inspecting from global context down to local details and deriving a final judgment, we design a dedicated CoT annotation format tailored to FAS. As shown in  Fig. 1, this format comprises six hierarchical levels: caption, facial description, facial attributes, reasoning, spoofing description, and conclusion. Second, we employ GPT-4o [27] to automatically generate initial CoT annotations. To guarantee data quality and accuracy, we adopt a human-in-the-loop workflow in which all automatically generated content is rigorously filtered and refined by expert annotators. The resulting high-quality subset is named FaceCoT-Gold100K. Third, to expand the dataset’s scale, we train a specialized FAS caption model on FaceCoT-Gold100K. During this process, we integrate a rule-verifiable Reinforcement Learning (RL) strategy to improve annotation quality and robustness across domains. The optimized model is capable of producing high-quality CoT annotations even on unseen data, allowing us to generate an additional 982K structured annotations, termed FaceCoT-Silver982K.
@@
To the best of our knowledge, FaceCoT is the first Visual Question Answering (VQA) dataset specifically designed for FAS, aggregating 1.08M training samples and covering 14 distinct attack types. The types of attacks included in the dataset are illustrated in  Fig. 2(a). Its carefully engineered, structured and hierarchical CoT annotation format not only ensures logical consistency and interpretability but also provides a natural learning pathway for downstream reasoning models. Moreover, we adopt a hybrid annotation workflow that combines GPT-4o–driven automatic generation with expert manual refinement, which further guarantees data quality and reliability. Finally, the integration of an RL–based strategy during the dataset expansion phase enhances cross-domain annotation accuracy and robustness, supplying the multimodal FAS community with unprecedentedly large-scale, high-quality training resources.
@@
The FaceCoT dataset provides a unique resource for training MLLMs. However, if trained in an end-to-end manner, the model is forced to learn CoT reasoning and classification at the same time, which leads to task interference and prevents the reasoning objective from fully converging. As a result, the model cannot fully exploit the fine-grained visual cues embedded in the CoT annotations. To address this, we propose a CoT-Enhanced Progressive Learning (CEPL) strategy, consisting of two stages: (1) Visual Enhancement Pre-training, we perform full-parameter Supervised Fine-Tuning (SFT) of the model using CoT data, thereby focusing the vision encoder on extracting fine-grained, spoof-relevant facial features; (2) Multi-task Joint Training, we inherit the vision encoder from the first stage, reset the connector and language decoder to their original pretrained weights, and fine-tune both the connector and decoder using LoRA modules. We then jointly train on CoT annotations and binary labels with a multi-task loss, ensuring that the model retains deep reasoning capabilities while rapidly adapting to the classification task.
@@
  • Proposing and releasing the FaceCoT dataset. To address the lack of VQA data in multimodal FAS, we introduce FaceCoT, the first VQA dataset designed specifically for FAS. It covers 14 attack types and comprises 1.08M samples. Each entry contains a structured visual CoT question-answering process, which guides models from image understanding to logical judgment, thereby improving detection accuracy and enhancing generalization ability.
  • Proposing a CoT-Enhanced Progressive Learning method. To fully leverage CoT data for FAS detection, we develop a CoT-Enhanced Progressive Learning approach that balances CoT reasoning and binary classification. This strategy significantly improves model performance on both reasoning and classification tasks.
  • Performance. Extensive experiments on FAS benchmarks demonstrate the value of our dataset and the effectiveness of our method. As shown in  Fig. 2(b), our approach outperforms state-of-the-art methods, achieving an average AUC improvement of 4.06% and an HTER reduction of 5.00%.
@@ 2 Related work
FAS is a key technology for protecting face recognition [8] systems from presentation attacks. FAS focuses on detecting these fraudulent attacks to ensure that the faces recognized by the system are genuine live faces. Early FAS methods primarily rely on handcrafted low-level feature extraction techniques, such as LBP [2], SIFT [28], and others. With advancements in deep learning, many FAS methods [24, 30] begin using models like CNNs and ViT to train classifiers. FAS methods [41, 44] gradually introduce multimodal learning, combining features from different types of data (e.g., RGB, IR, depth) to improve model performance. These methods achieve good results in intra-dataset scenarios but struggle with generalization to unseen attack types from out-of-domain data. As a result, Domain Adaptation (DA) [7, 4] and Domain Generalization (DG) [7, 11] approaches are developed. DA minimizes the distribution gap between the source and target domains by utilizing unlabeled target data. However, in practice, collecting unlabeled target data is challenging. DG learns broadly applicable features from multiple source domains, enabling good predictions in the target domain. However, due to the diversity of attack types and data collection methods, it is difficult to find a universal feature space for fake faces, leading to insufficient generalization. Additionally, these methods lack interpretability, making it challenging to understand the decision-making process of the model, which is crucial in high-stakes applications like security systems.
@@
In MLLMs, CoT is a method for solving complex problems through a series of intermediate reasoning steps, providing an interpretable reasoning path for the model when answering questions. Recent research [37, 35] has shown that the CoT prompting method significantly enhances the reasoning abilities of large language models in reasoning tasks. Consequently, researchers have attempted to apply CoT in MLLMs. Some researchers define their own reasoning stages. For example, LLaVA-CoT [39] proposes a method combining CoT with MLLMs by designing a ‘summary-caption-reasoning-conclusion’ reasoning process to enhance the model’s reasoning ability. Other researchers have explored allowing models to autonomously design the reasoning stages, such as PS-CoT [17], which enables large language models to generate task-solving plans before generating reasoning evidence. CoT-PT [9] adopts a hierarchical reasoning approach, moving from abstract concepts to specific ones. BDoG [47] employs a unique approach using three agents who repeatedly debate to implicitly form a reasoning graph that explores and aggregates various thoughts. However, in the field of FAS, such advanced CoT-based reasoning strategies have not yet been fully explored or applied.
@@ 3 The FaceCoT Dataset
In this section, we introduce the FaceCoT dataset, comprising FaceCoT-Gold100K and FaceCoT-Silver982K, and detail the entire construction process. Figure 3 provides an overview of the pipeline. Subsection 3.1 describes how we collect and curate our raw data to form FaceCoT-Gold100K. Subsection 3.2 provides a detailed explanation of how we design the CoT structure for FAS and use GPT-4o for data annotation. Subsection 3.3 presents the training of our FAS caption model, which is based on FaceCoT-Gold100K under an SFT augmented by RL, and how we then apply it to annotate the complete raw dataset, yielding the larger FaceCoT-Silver982K dataset.
@@ 3.1 Data collection
To construct the FaceCoT dataset, our objective is to gather images with large scale, broad attack coverage, and strong demographic diversity to support reliable FAS. We therefore select images that represent both genuine faces and diverse spoofing types, ensuring inclusion of mainstream 2D and 3D presentation attacks under varied conditions. Following these criteria, we collect images from two major FAS datasets: CelebA-Spoof [45], which provides 625K images from 10K subjects with genuine faces and 10 spoofing types, and Wild-FAS (WFAS) [34], which contributes 1.38M images (853K spoofed and 529K genuine) from a large number of subjects and 14 attack types captured in unconstrained real-world environments. This combination enables FaceCoT to achieve both high diversity and strong realism.
@@ 3.2 Data annotation
Humans typically judge authenticity of images by following a hierarchical “global-to-local” reasoning path: they first assess the overall scene, then focus on facial details and attributes, and finally draw a conclusion through logical analysis. To emulate this process, we partition our CoT annotations into six modules, enabling the model to perform human-like detection reasoning:
@@
    Facial Description: This part focuses the model’s attention on the facial region and requires a description of the facial features. The face is the most easily simulated area in spoofing attacks, so this section helps the model concentrate on potential spoofing regions.
@@
    Facial Attributes: Further, the model is asked to describe various facial attributes, including facial features, textures, and expressions. By describing these attributes, the model’s ability to perceive fine-grained facial details is enhanced.
@@
    Conclusion: This section is the final summary of all the reasoning conducted previously.
Based on the above design approach, we further standardize the annotation format of FaceCoT data, adopting the following structure: <Caption></Caption>, <Facial Description></Facial Description>, <Facial Attributes></Facial Attributes>, <Reasoning></Reasoning>, <Spoofing Descript- ion></Spoofing Description>, <Conclusion></ Conclusion>, thus providing the model with clear and structured input, which helps improve the accuracy and stability of multimodal reasoning (as shown in  Fig. 1).
@@
After designing the CoT format for the answers, we employ GPT-4o [27] to conduct the annotation process. During annotation, for different spoofing types, we provide corresponding hints (e.g., classification standards such as ”photographing a poster constitutes spoofing.”) to prevent the model from identifying features but failing to properly control the decision boundaries.
@@
After completing the initial annotation, we compare the extracted labels (obtained via regular expressions) with the original ground-truth labels. A total of 98,976 samples are correctly annotated. For those incorrectly annotated, we perform a second round of annotation using GPT-4o. Samples that remain incorrectly labeled after this second attempt are marked as hard cases, resulting in 581 such instances. For these hard cases, professional annotators perform cleaning. The annotators review the original labels, locate the corresponding stages, and correct any deficiencies. Finally, they reassemble a logically coherent CoT text. The detailed hard case cleaning process is provided in the Appendix.
@@ 3.3 Data expansion
Although we assembled a diverse FaceCoT-Gold100K, it still falls short of covering every attack type in real-world FAS scenarios. Expanding the dataset with GPT-4o plus human annotation would be costly and hard to scale. Therefore, we train a caption model to address this issue. We first train a caption model on FaceCoT-Gold100K via SFT, but find that its auto-generated CoT annotations on unseen images suffer from two issues: semantic errors and format errors. To fix this, we adopt a verified reinforcement fine-tuning scheme inspired by VRFT [26], designing rewards for both accuracy and format compliance. Specifically, an accuracy reward gives 1 if <Conclusion> matches the image’s ground-truth label and 0 otherwise, while a format reward checks whether the output follows the FaceCoT template. This RL step markedly improves annotation correctness and offers a scalable path to enlarging FaceCoT. The details of the RL procedure are provided in the Appendix.
@@
We apply the trained caption model to annotate the training splits of CelebA-Spoof and WFAS. Annotation accuracy is measured by (1) regular-expression checks for template compliance and (2) consistency between the generated <Conclusion> tag and the ground-truth label. With standard SFT, we achieve an accuracy of 88%, whereas reinforcement fine-tuning raises this figure to 99.6%. We ultimately create FaceCoT-Silver982K, an additional 982K high-quality FAS CoT annotations.